OpenStudy (jjwalton04):

Find the ordered pairs for the x- and y-intercepts of the equation 5x − 6y = 30 and select the appropriate option below

OpenStudy (jjwalton04):

The x-intercept is (6, 0), the y-intercept is (0, –5). The x-intercept is (0, 6); the y-intercept is (–5, 0). The x-intercept is (5, 0); the y-intercept is (0, –6). The x-intercept is (0, 5); the y-intercept is (–6, 0).

pooja195 (pooja195):

To find the x intercept: 5x − 6y = 30 5x − 6(0) = 30 <---set y value as 0 5x=30 <---divide by 5 on both sides x=? To find y intercept 5x − 6y = 30 5(0) − 6y = 30 <---set x value as 0 -6y=30 <---divide both sides by -6 y=?

OpenStudy (jjwalton04):

so y=5 and x=6

pooja195 (pooja195):

y=-5 x=6 is correct good job! :) Welcome to openstudy! :D

OpenStudy (jjwalton04):

lolz thanks

OpenStudy (jjwalton04):

btw @pooja195 is it (5,0) or (0,5)

pooja195 (pooja195):

(x,y) (x,0) (0,y)